# Exclusive $J/\psi$, $\psi(2s)$, and $e^{+}e^{-}$ pair production in Au$+$Au ultra-peripheral collisions at RHIC

### v1

@@ V.1 Momentum transfer distributions
The square of the four-momentum transferred, $-t$, between the incoming and outgoing nucleus (transferred along the gluon lines in Fig. 1) characterizes the underlying nuclear geometry and its fluctuation in photon-nucleus interactions. It has both longitudinal and transverse components:
@@
$-t_{\parallel}=M^{2}_{J/\psi}/(\gamma^{2}e^{\pm y}),$ (5)
$-t_{\perp}=p^{2}_{\rm{T}}.$ (6)
@@
where $p_{\rm{T}}$ is the transverse momentum of the $J/\psi~{}$. The measured $p^{2}_{\rm{T}}$ of the $J/\psi~{}$deviates slightly from $-t$ due to the transverse momentum of the hard photon emitted by the other nucleus, on average $\sim 30$ MeV/c. However, this smearing due to the photon $p_{\rm{T}}$ is relatively small and is included in some model calculations [27, 20, 2], which can be directly compared to the measured $p^{2}_{\rm{T}}$ distributions of the $J/\psi~{}$in the data.
@@
Figure 5: Differential cross section $d^{2}\sigma/dp^{2}_{\mathrm{T}}dy$ of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$($0<p^{2}_{\rm{T}}<0.12~{}\rm(GeV/c)^{2}$) in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm{NN}}}}=200$ GeV. The rapidity of the $J/\psi~{}$is $|y|<1.0$ and averaged $W_{\mathrm{\gamma*N}}~{}$is 25.0 GeV. MC model STARlight (upper left), nuclear shadowing of leading twist approximation calculation (upper right), Sartre MC prediction (lower left), and the color-glass-condensate prediction (lower right) are compared with the data, presented as lines. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. The ratio between data and models is shown in the lower sub-panel of each figure. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown.
@@
In Fig. 4 and Fig. 5, the $J/\psi~{}$differential cross section, $d^{2}\sigma/dp^{2}_{\rm{T}}dy$ over the full measured range of $p^{2}_{\rm{T}}<2.2$ $\rm(GeV/c)^{2}$ and the low momentum transfer range $p^{2}_{\rm{T}}<0.12$ $\rm(GeV/c)^{2}$ is shown, respectively. The data are identical among the four panels, where each panel shows a comparison to a different theoretical model, indicated by the legend. At lowest $p^{2}_{\rm{T}}$, the data exhibit a steep peak characteristic of coherent $J/\psi~{}$photoproduction. As $p^{2}_{\rm{T}}$ increases the data follow a softly falling exponential, indicative of incoherent photoproduction with scattering off individual nucleons in the nucleus. At higher $p^{2}_{\rm{T}}$, the distribution flattens in the region where the scattered nucleon dissociates. The models compared with the data are the STARlight event generator, nuclear shadowing model LTA with weak shadowing mode, Sartre model with sub-nucleonic fluctuations, and CGC calculation with sub-nucleonic fluctuations. All models exhibit similar feature as the data, while the second peak or shoulder structure is caused by the coherent contribution. Furthermore, all models are presented with coherent, incoherent, and total cross sections in Fig. 5, while only the total cross sections are shown in Fig. 4. Ratios of data to the model total cross sections are shown at the bottom of each panel, where the model values are calculated by the integral of each bin instead of at the bin center.
@@
Based on the data and model comparisons, the prediction of Sartre and LTA are found to be better in describing the magnitudes and slopes of the coherent and incoherent components. The Sartre model has the same underlying physics model as the CGC framework with minor differences in the implementation. The CGC additionally accounts for the initial photon transverse momentum and interference effects. Both models consider the i) nonlinear gluon evolution in the target nucleus and ii) sub-nucleonic gluon density fluctuations. The CGC calculation overpredicts the cross section by a factor of $\sim$1.5 at low ($\sim 0.01~{}\rm{(GeV/c)^{2}}$) and high ($>1~{}\rm{(GeV/c)^{2}}$) $p^{2}_{\mathrm{T}}~{}$when compared to the data. The same overprediction factor has been found in comparison to the LHC data [27]. Note that this prediction is explicitly made for the RHIC UPC measurement.
@@
Figure 6 shows the $p^{2}_{\rm{T}}$ distributions for all four neutron categories over the full rapidity range $|y|<1$. The distinction between the coherent peak and incoherent tail is clearly visible due to the large difference in the slope, suggesting that the total cross section for each process can be measured with minimal extrapolation. The template fits to the cross section for coherent and incoherent $J/\psi~{}$production and their sum are also shown in Fig. 6. The coherent cross section is determined by integrating the data over $p^{2}_{\rm{T}}<0.09$ $\rm(GeV/c)^{2}$ and subtracting the incoherent template over the same range. The incoherent cross section is obtained by integrating the data over $p^{2}_{\rm{T}}>0.025$ $\rm(GeV/c)^{2}$ and subtracting the coherent template over the same range. To account for the low $p^{2}_{\mathrm{T}}~{}$region, the cross section is scaled by the ratio of the incoherent template integrated over all $p^{2}_{\rm{T}}$ to that of $p^{2}_{\rm{T}}>0.025$ $\rm(GeV/c)^{2}$. These cross sections were measured for each of the three rapidity bins, resulting in the differential cross sections $d\sigma/dy$.
@@ V.2 Rapidity distributions
In Fig. 7, the differential cross section, $d\sigma/dy$, of $J/\psi~{}$photoproduction in the coherent (top left) and incoherent (top right) processes are presented. The incoherent to coherent ratio is shown in the bottom panel. The measurements are also separated in different neutron emission configurations. The CGC calculation predicts that the ratio between the rapidity-dependent cross sections for incoherent and coherent $J/\psi~{}$photoproduction is 40%, which is supported by the STAR data shown in Fig. 7, as opposed to 20% [27]. Although this is model dependent, the enhancement of incoherent cross section has been regarded as a result of the sub-nucleonic parton density fluctuation observed in the HERA data [54].
@@ V.3 Coherent $J/\psi~{}$cross section in $\gamma$+Au collisions
Figure 8: Total coherent $J/\psi~{}$photoproduction cross section as a function of $W_{\mathrm{\gamma*N}}~{}$in Au$+$Au UPCs. The data are compared with an expectation of a free nucleon provided by the Impulse Approximation (IA) [55] and color glass condensate (CGC) [27]. The ratio between data and the Impulse Approximation at $W_{\mathrm{\gamma*N}}~{}$$=25.0$ GeV is the suppression factor, shown in the figure. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown on the data points.
@@
With the differential cross sections $d\sigma/dy$ for the three neutron categories $0n0n$, $0nXn$, and $XnXn$, the procedure to extract the $\gamma$+Au cross section described in Section IV.5 was followed. Each rapidity measurement determines two cross sections at different photon-nucleus center-of-mass energies $W_{\pm}$. For the present data the results for the higher energy $W_{+}$ are not statistically significant (consistent with zero) due to the low fraction of high energy flux compared to low energy. For the lowest rapidity interval the lower energy result at $W_{-}$ is also not significant. Instead, the rapidity range $|y|<0.2$ was used to measure the $\gamma$+Au cross section at the mean $W$ = 25.0 GeV. Note that there is a --0.5 GeV shift in the estimate of $W_{\mathrm{\gamma*N}}~{}$for rapidity $|y|<0.2$, caused by the higher photon flux of the lower energy photon contribution; however, the effect of this shift is found to be negligible.
@@
In Fig. 8, the total coherent $J/\psi~{}$photoproduction as a function of $W_{\mathrm{\gamma*N}}~{}$is presented. The data are free of photon energy ambiguity, based on the method discussed in Sec. IV.5. The data are found to be suppressed with respect to the IA[55] for all measured energies. Quantitatively, the suppression factor $\rm S^{Au}_{coh}$ is reported, which is the ratio between coherent $J/\psi~{}$cross section and the IA. It is found that $\rm S^{Au}_{coh}=0.71\pm 0.04\pm 0.07\pm 0.07$ at $W_{\mathrm{\gamma*N}}~{}$$=25.0$ GeV. The first quoted error is the model uncertainty on IA [55] for Au nucleus and the second error is a combination of statistics and systematic uncertainties added in quadrature, while the third is from the scale uncertainty of the integrated luminosity. The reported STAR results in this analysis are the first measurements that contain no photon energy ambiguity at $y\neq 0$ in a symmetric collision system. The data can only be compared with the nuclear shadowing model, since the saturation-based models do not apply to $W_{\mathrm{\gamma*N}}~{}$lower than $\sim 31$ GeV. It is found that the LTA with weak shadowing gives an excellent description of the data, while the LTA with strong shadowing predicts a slighly lower cross section. The precise data presented here provide a stringent constraint on the nuclear PDFs and important information towards the understanding of the fundamental mechanism of such parton modification.
@@ V.4 Incoherent $J/\psi~{}$cross section
In Fig. 9, the differential cross section of $J/\psi~{}$production as a function of $p^{2}_{\mathrm{T}}~{}$is shown for $|y|<1.0$ for the $all~{}n$ neutron class. Due to the large rapidity range, the $\langle W_{\gamma^{\ast}N}\rangle$ is estimated to be 19.0 GeV based on the $all~{}n$ photon flux [2] and the energy-dependent cross section in $ep$ photoproduction [51].
Starting above $p^{2}_{\mathrm{T}}~{}$$>0.02\rm~{}(GeV/c)^{2}$, the cross section is dominated by the incoherent production. In order to compare with a free proton, the H1 published fit [51] to $ep$ collisions is scaled down from $\langle W_{\gamma^{\ast}N}\rangle$ = 55 GeV to 19.0 GeV, based on the well measured energy-dependent cross sections parametrization as follows [51]:
$\sigma_{el}=N_{el}\cdot(W_{\rm{\gamma*N}}/90)^{\delta_{el}},$ (8)
$\sigma_{pd}=N_{pd}\cdot(W_{\rm{\gamma*N}}/90)^{\delta_{pd}}.$ (9)
Here the $\sigma_{el}$ and $\sigma_{pd}$ are the proton elastic and proton dissociation cross section as a function of $W_{\mathrm{\gamma*N}}$. The parameters are $N_{el}=81\pm 3\rm{nb}$, $\delta_{el}=0.67\pm 0.03$, $N_{pd}=66\pm 7\rm{nb}$, and $\delta_{pd}=0.42\pm 0.05$. For the elastic proton case, UPC measurements in proton-lead UPCs at the LHC [56] has a similar parametrization. Thus, the cross section ratios at different energies (e.g., 19.0 and 25.0 GeV) with respect to the H1 measured energy (55 GeV) are derived, which are used for obtaining the differential cross section $d\sigma/dt$ for proton elastic and proton dissociation at the STAR UPC kinematics.
@@
$d^{2}\sigma_{\rm Au+Au\rightarrow J/\psi+Y}(\langle W_{\gamma^{\ast}N}\rangle)/dp^{2}_{\mathrm{T}}dy=$
$2\Phi^{\rm all~{}n}_{T,\gamma}(\langle W_{\gamma^{\ast}p}\rangle)\cdot A\cdot\left[d\sigma_{\rm{\gamma^{\ast}+p\rightarrow J/\psi+Y}}(\langle W_{\gamma^{\ast}p}\rangle)/dt\right].$ (12)
Here the $A$ is 197 for the Au nucleus, $\Phi^{\rm all~{}n}_{T,\gamma}=5.03$ is the average transverse photon flux at $y=0$ (the coefficient $2$ is the total flux from both beams being photon emitters), and $d\sigma_{\rm{\gamma^{\ast}+p\rightarrow J/\psi+Y}}(\langle W_{\gamma^{\ast}p}\rangle)/dt$ is the published H1 data at $\langle W_{\rm\gamma^{\ast}N}\rangle=\langle W_{\rm\gamma^{\ast}p}\rangle=19.0$ GeV scaled down from 55 GeV for both the elastic proton ($\rm{Y}=p$) and the proton dissociation ($\rm{Y}\neq p$). The notation is similar for the STAR data, where $\rm{Y}$ can be elastic nucleon or nucleon dissociation. Note that the published H1 data had been corrected for photon flux that is integrated over the phase space of $W_{\rm\gamma^{\ast}p}$, which is equivalent to the normalization of $1/dy$ in UPC measurements. The equivalent Au$+$Au UPC cross section for the free proton data is shown as the black solid line in Fig 9, where the uncertainty band is propagated from the errors of the parametrization.
Moreover, we use the H1 free proton data as a template to fit the STAR data with only the normalization constant as a free parameter. The integral of $d^{2}\sigma/dp^{2}_{\rm{T}}dy$ from $p^{2}_{\mathrm{T}}~{}$= 0 to 2.2 $\rm{(GeV/c)^{2}}$ between the fit and the H1 data is defined as the incoherent suppression factor, $\rm S^{Au}_{incoh}$. It is found that the $\rm S^{Au}_{incoh}$ is $0.49^{+0.04}_{-0.05}\pm 0.05\pm 0.05$ at $W_{\mathrm{\gamma*N}}~{}$$=19.0~{}\rm{GeV}$. For $W_{\mathrm{\gamma*N}}~{}$$=25.0~{}\rm{GeV}$ corresponding to the measurement within rapidity range $|y|<0.2$, the same procedure has been performed and the suppression factor is found to be $0.36^{+0.03}_{-0.04}\pm 0.04\pm 0.04$. Here the first uncertainty is the H1 parametrization uncertainty [51], the second one is from the measurement that includes statistical and systematic uncertainty, and the third is the scale uncertainty on the integrated luminosity. Therefore, the nuclear suppression in incoherent $J/\psi~{}$photoproduction in Au$+$Au UPCs has been found to be stronger than that in the coherent case. This has been qualitatively predicted by the nuclear shadowing model LTA [17, 18].
@@
Furthermore, the data are compared with different models. For the CGC calculations, the data are found to be in between the scenarios of strong sub-nucleonic parton density fluctuations and no fluctuations. It is not clear that the data directly supports either scenario. For the Sartre model, similar sub-nucleonic parton density fluctuations are included, which describes well the high $p^{2}_{\mathrm{T}}~{}$tail but not the low $p^{2}_{\mathrm{T}}~{}$behavior. For the LTA with weak shadowing, the description of the data is very good. However, this is expected as the LTA model uses the HERA data parametrization. Finally, for the BeAGLE event generator, the cross section is underestimated for the entire $p^{2}_{\mathrm{T}}~{}$range, which indicates using only the nuclear PDF, e.g. EPS09 [57], is not sufficient to describe the data.

### v3

@@ V.1 Momentum transfer distributions
The square of the four-momentum transferred, −$-t$, between the incoming and outgoing nucleus (transferred along the gluon lines in Fig. 1) characterizes the underlying nuclear geometry and its fluctuation in photon-nucleus interactions. It has both longitudinal and transverse components:
@@
−$-t_{\parallel}=M^{2}_{J/\psi}/(\gamma^{2}e^{\pm y}),$ (5)
−$-t_{\perp}=p^{2}_{\rm{T}}.$ (6)
@@
where $p_{\rm{T}}$ is the transverse momentum of the $J/\psi~{}$. The measured $p^{2}_{\rm{T}}$ of the $J/\psi~{}$deviates slightly from −$-t$ due to the transverse momentum of the hard photon emitted by the other nucleus, on average ≈$\approx 30$ MeV/c. However, this smearing due to the photon $p_{\rm{T}}$ is relatively small and is included in some model calculations [28, 20, 2], which can be directly compared to the measured $p^{2}_{\rm{T}}$ distributions of the $J/\psi~{}$in the data.
@@
Figure 5: Differential cross section $d^{2}\sigma/dp^{2}_{\mathrm{T}}dy$ of $J/\psi~{}$photoproduction as a function of $p^{2}_{\mathrm{T}}~{}$($0<p^{2}_{\rm{T}}<0.12~{}\rm[(GeV/c)^{2}]$) in Au$+$Au UPCs at $\sqrt{s_{{}_{\rm{NN}}}}=200$ GeV. The rapidity of the $J/\psi~{}$is $|y|<1.0$ and averaged $W_{\mathrm{\gamma N}}~{}$is 25.0 GeV. MC model STARlight (upper left), nuclear shadowing of leading twist approximation calculation (upper right), Sartre MC prediction (lower left), and the color-glass-condensate prediction (lower right) are compared with the data, presented as lines. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. The ratio between data and models is shown in the lower sub-panel of each figure. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown.
@@
In Fig. 4 and Fig. 5, the $J/\psi~{}$differential cross section, $d^{2}\sigma/dp^{2}_{\rm{T}}dy$ over the full measured range of $p^{2}_{\rm{T}}<2.2$ $\rm[(GeV/c)^{2}]$ and the low momentum transfer range $p^{2}_{\rm{T}}<0.12$ $\rm[(GeV/c)^{2}]$ is shown, respectively. The data are identical among the four panels, where each panel shows a comparison to a different theoretical model, indicated by the legend. At lowest $p^{2}_{\rm{T}}$, the data exhibit a steep peak characteristic of coherent $J/\psi~{}$photoproduction. As $p^{2}_{\rm{T}}$ increases the data follow a softly falling exponential, indicative of incoherent photoproduction with scattering off individual nucleons in the nucleus. At higher $p^{2}_{\rm{T}}$, the distribution flattens in the region where the scattered nucleon dissociates. The models compared with the data are the STARlight event generator, nuclear shadowing model LTA with weak shadowing mode, Sartre model with sub-nucleonic fluctuations, and CGC calculation with sub-nucleonic fluctuations. All models exhibit similar feature as the data, while the second peak or shoulder structure is caused by the coherent contribution. Furthermore, all models are presented with coherent, incoherent, and total cross sections in Fig. 5, while only the total cross sections are shown in Fig. 4. Ratios of data to the model total cross sections are shown at the bottom of each panel, where the model values are calculated by the integral of each bin instead of at the bin center.
@@
Based on the data and model comparisons, the prediction of Sartre and LTA are found to be better in describing the magnitudes and slopes of the coherent and incoherent components. The Sartre model has the same underlying physics model as the CGC framework with minor differences in the implementation. The CGC additionally accounts for the initial photon transverse momentum and interference effects. Both models consider the i) nonlinear gluon evolution in the target nucleus and ii) sub-nucleonic gluon density fluctuations. The CGC calculation overpredicts the cross section by a factor of $\approx$1.5 at low (≈$\approx 0.01~{}\rm{[(GeV/c)^{2}]}$) and high ($>1~{}\rm{[(GeV/c)^{2}]}$) $p^{2}_{\mathrm{T}}~{}$when compared to the data. The same overprediction factor has been found in comparison to the LHC data [28]. Note that this prediction is explicitly made for the RHIC UPC measurement.
@@
Figure 6 shows the $p^{2}_{\rm{T}}$ distributions for all four neutron categories over the full rapidity range $|y|<1$. The distinction between the coherent peak and incoherent tail is clearly visible due to the large difference in the slope, suggesting that the total cross section for each process can be measured with minimal extrapolation. The template fits to the cross section for coherent and incoherent $J/\psi~{}$production and their sum are also shown in Fig. 6. The coherent cross section is determined by integrating the data over $p^{2}_{\rm{T}}<0.09$ $\rm[(GeV/c)^{2}]$ and subtracting the incoherent template over the same range. The incoherent cross section is obtained by integrating the data over $p^{2}_{\rm{T}}>0.025$ $\rm[(GeV/c)^{2}]$ and subtracting the coherent template over the same range. To account for the low $p^{2}_{\mathrm{T}}~{}$region, the cross section is scaled by the ratio of the incoherent template integrated over all $p^{2}_{\rm{T}}$ to that of $p^{2}_{\rm{T}}>0.025$ $\rm[(GeV/c)^{2}]$. These cross sections were measured for each of the three rapidity bins, resulting in the differential cross sections $d\sigma/dy$.
@@ V.2 Rapidity distributions
In Fig. 7, the differential cross section, $d\sigma/dy$, of $J/\psi~{}$photoproduction in the coherent (top left) and incoherent (top right) processes are presented. The incoherent to coherent ratio is shown in the bottom panel. The measurements are also separated in different neutron emission configurations. The CGC calculation predicts that the ratio between the rapidity-dependent cross sections for incoherent and coherent $J/\psi~{}$photoproduction is 40%, for nucleons that have event-by-event fluctuation in their gluon density [28]. On the other hand, with a static nucleon without the fluctuation, the prediction is around 20%. The STAR data is found to be close to 40%, which is shown in Fig. 7, and favors the case of a fluctuating nucleon on this particular observable [28]. Although this is model dependent, the enhancement of incoherent cross section has been regarded as a result of the sub-nucleonic parton density fluctuation observed in the HERA data [56].
@@ V.3 Coherent $J/\psi~{}$cross section in $\gamma$+Au collisions
Figure 8: Total coherent $J/\psi~{}$photoproduction cross section as a function of $W_{\mathrm{\gamma N}}~{}$in Au$+$Au UPCs. The data are compared with an expectation of a free nucleon provided by the Impulse Approximation (IA) [57] and color glass condensate (CGC) [28]. The ratio between data and the Impulse Approximation at $W_{\mathrm{\gamma N}}~{}$$=25.0$ GeV is the suppression factor, shown in the figure. Statistical uncertainty is represented by the error bars, and the systematic uncertainty is denoted as boxes. There is a systematic uncertainty of 10% from the integrated luminosity that is not shown on the data points.
@@
With the differential cross sections $d\sigma/dy$ for the three neutron categories $0n0n$, $0nXn$, and $XnXn$, the procedure to extract the $\gamma$+Au cross section described in Section IV.5 was followed. Each rapidity measurement determines two cross sections at different photon-nucleus center-of-mass energies $W_{\pm}$. For the present data the results for the higher energy $W_{+}$ are not statistically significant (consistent with zero) due to the low fraction of high energy flux compared to low energy. For the lowest rapidity interval the lower energy result at $W_{-}$ is also not significant. Instead, the rapidity range $|y|<0.2$ was used to measure the $\gamma$+Au cross section at the mean $W$ = 25.0 GeV. Note that there is a −--0.5 GeV shift in the estimate of $W_{\mathrm{\gamma N}}~{}$for rapidity $|y|<0.2$, caused by the higher photon flux of the lower energy photon contribution; however, the effect of this shift is found to be negligible.
@@
In Fig. 8, the total coherent $J/\psi~{}$photoproduction as a function of $W_{\mathrm{\gamma N}}~{}$is presented. The data are free of photon energy ambiguity, based on the method discussed in Sec. IV.5. The data are found to be suppressed with respect to the IA[57] for all measured energies. Quantitatively, the suppression factor $\rm S^{Au}_{coh}$ is reported, which is the ratio between coherent $J/\psi~{}$cross section and the IA. It is found that $\rm S^{Au}_{coh}=0.71\pm 0.04\pm 0.07\pm 0.07$ at $W_{\mathrm{\gamma N}}~{}$$=25.0$ GeV. The first quoted error is the model uncertainty on IA [57] for Au nucleus and the second error is a combination of statistics and systematic uncertainties added in quadrature, while the third is from the scale uncertainty of the integrated luminosity. The reported STAR results in this analysis are the first measurements that contain no photon energy ambiguity at $y\neq 0$ in a symmetric collision system at RHIC. The data can only be compared with the nuclear shadowing model, since the saturation-based models do not apply to $W_{\mathrm{\gamma N}}~{}$lower than ≈$\approx 31$ GeV. It is found that the LTA with weak shadowing gives an excellent description of the data, while the LTA with strong shadowing predicts a slighly lower cross section. The data are found to be consistent with the LHC data at a similar energy range [11, 10]. The implication of nuclear shadowing and its impact on the nuclear PDFs should be further investigated within its theoretical framework, and future p+A data with hard diffraction may shine new light on this question. Nevertheless, the precise data presented here provide a stringent constraint on the nuclear PDFs and important information towards the understanding of the fundamental mechanism of such parton modification.
@@ V.4 Incoherent $J/\psi~{}$cross section
In Fig. 9, the differential cross section of $J/\psi~{}$production as a function of $p^{2}_{\mathrm{T}}~{}$is shown for $|y|<1.0$ for the $all~{}n$ neutron class. Due to the large rapidity range, the ⟨$\langle W_{\gamma N}\rangle$ is estimated to be 19.0 GeV based on the $all~{}n$ photon flux [2] and the energy-dependent cross section in $ep$ photoproduction [52].
Starting above $p^{2}_{\mathrm{T}}~{}$$>0.02\rm~{}[(GeV/c)^{2}]$, the cross section is dominated by the incoherent production. In order to compare with a free proton, the H1 published fit [52] to $ep$ collisions is scaled down from ⟨$\langle W_{\gamma N}\rangle$ = 55 GeV to 19.0 GeV, based on the well measured energy-dependent cross sections parametrization as follows [52]:
$\sigma_{el}=N_{el}\cdot(W_{\rm{\gamma N}}/90)^{\delta_{el}},$ (8)
$\sigma_{pd}=N_{pd}\cdot(W_{\rm{\gamma N}}/90)^{\delta_{pd}}.$ (9)
Here the $\sigma_{el}$ and $\sigma_{pd}$ are the proton elastic and proton dissociation cross section as a function of $W_{\mathrm{\gamma N}}$. The parameters are $N_{el}=81\pm 3\rm{nb}$, $\delta_{el}=0.67\pm 0.03$, $N_{pd}=66\pm 7\rm{nb}$, and $\delta_{pd}=0.42\pm 0.05$. For the elastic proton case, UPC measurements in proton-lead UPCs at the LHC [58] has a similar parametrization. Thus, the cross section ratios at different energies (e.g., 19.0 and 25.0 GeV) with respect to the H1 measured energy (55 GeV) are derived, which are used for obtaining the differential cross section $d\sigma/dt$ for proton elastic and proton dissociation at the STAR UPC kinematics.
@@
$d^{2}\sigma_{\rm Au+Au\rightarrow J/\psi+Y}(\langle W_{\gamma N}\rangle)/dp^{2}_{\mathrm{T}}dy=$
$2\Phi^{\rm all~{}n}_{T,\gamma}(\langle W_{\gamma p}\rangle)\cdot A\cdot\left[d\sigma_{\rm{\gamma+p\rightarrow J/\psi+Y}}(\langle W_{\gamma p}\rangle)/dt\right].$ (12)
Here the $A$ is 197 for the Au nucleus, $\Phi^{\rm all~{}n}_{T,\gamma}=5.03$ is the average transverse photon flux at $y=0$ (the coefficient $2$ is the total flux from both beams being photon emitters), and $d\sigma_{\rm{\gamma+p\rightarrow J/\psi+Y}}(\langle W_{\gamma p}\rangle)/dt$ is the published H1 data at ⟨$\langle W_{\rm\gamma N}\rangle=\langle W_{\rm\gamma p}\rangle=19.0$ GeV scaled down from 55 GeV for both the elastic proton ($\rm{Y}=p$) and the proton dissociation ($\rm{Y}\neq p$). The notation is similar for the STAR data, where $\rm{Y}$ can be elastic nucleon or nucleon dissociation. Note that the published H1 data had been corrected for photon flux that is integrated over the phase space of $W_{\rm\gamma p}$, which is equivalent to the normalization of $1/dy$ in UPC measurements. The equivalent Au$+$Au UPC cross section for the free proton data is shown as the black solid line in Fig 9, where the uncertainty band is propagated from the errors of the parametrization.
Moreover, we use the H1 free proton data as a template to fit the STAR data with only the normalization constant as a free parameter. The integral of $d^{2}\sigma/dp^{2}_{\rm{T}}dy$ from $p^{2}_{\mathrm{T}}~{}$= 0 to 2.2 $\rm{[(GeV/c)^{2}]}$ between the fit and the H1 data is defined as the incoherent suppression factor, $\rm S^{Au}_{incoh}$. It is found that the $\rm S^{Au}_{incoh}$ is $0.49^{+0.04}_{-0.05}\pm 0.05\pm 0.05$ at $W_{\mathrm{\gamma N}}~{}$$=19.0~{}\rm{GeV}$. For $W_{\mathrm{\gamma N}}~{}$$=25.0~{}\rm{GeV}$ corresponding to the measurement within rapidity range $|y|<0.2$, the same procedure has been performed and the suppression factor is found to be $0.36^{+0.03}_{-0.04}\pm 0.04\pm 0.04$. Here the first uncertainty is the H1 parametrization uncertainty [52], the second one is from the measurement that includes statistical and systematic uncertainty, and the third is the scale uncertainty on the integrated luminosity. Therefore, the nuclear suppression in incoherent $J/\psi~{}$photoproduction in Au$+$Au UPCs has been found to be stronger than that in the coherent case. This has been qualitatively predicted by the nuclear shadowing model LTA [17, 18].
@@
Furthermore, the data are compared with different models. For the CGC calculations, the data are found to be in between the scenarios of strong sub-nucleonic parton density fluctuations and no fluctuations. It is not clear that the data directly supports either scenario. For the Sartre model, similar sub-nucleonic parton density fluctuations are included, which describes well the high $p^{2}_{\mathrm{T}}~{}$tail but not the low $p^{2}_{\mathrm{T}}~{}$behavior. Note that both the CGC and the Sartre model are calculated based on a higher energy configuration (corresponding to $x=0.01$) due to their model limitations. For the LTA with weak shadowing, the description of the data is very good. However, this is expected as the LTA model uses the HERA data parametrization. Finally, for the BeAGLE event generator, the cross section is underestimated for the entire $p^{2}_{\mathrm{T}}~{}$range, which indicates using only the nuclear PDF, e.g. EPS09 [59], is not sufficient to describe the data.